# DDN-SLAM: Real-time Dense Dynamic Neural Implicit SLAM with Joint Semantic Encoding

### 03 Jan 2024 v1

@@ 3.2 Volume Rendering and Semantic Encoding
For representing 3D scenes in the presence of occlusion, we employ a volumetric rendering process combined with multi-resolution hash encoding [25]. This allows for effective geometric detail representation and adaptability to large-scale scenes[83], where each corner of a unit cell corresponds to a hashed grid that embeds encoding features. The given position $x$ can be mapped to the respective positions on each resolution grid, and trilinear interpolation is performed on the grid cell corners to obtain concatenated encoding feature vectors $\boldsymbol{\gamma}_{t}(\mathbf{x}_{t},\mathbf{t})\in\mathbb{R}^{c}$ After feature concatenation, we input them into an MLP. More specifically, our SDF network $F_{\Phi_{sdf}}$ is composed of a single multi-layer perceptron (MLP), with learnable parameters $\Phi_{sdf}$. It takes the point position $x$ and the corresponding hash encoding $h_{\Phi_{hash}}(\mathbf{x})$ as inputs and predicts the signed distance field (SDF) as:
@@
where $\Omega$ is the sigmoid function. $T_{k}=\prod_{j=1}^{k-1}\left(1-\alpha_{j}\right)$ denotes the cumulative transmission rate, which represents the proportion of light reaching the camera. Similar to [25], we utilize skip voxel strategy. We maintain a series of K multi-scale occupancy grids. The occupancy status of each grid is continuously traversed using a Digital Differential Analyzer (DDA)[12] after 16 training iterations, and grid updates are performed. To improve efficiency, we employ a ray propagation sampling method to reduce unnecessary sampling. When $T_{k}$ drops below a threshold $\varepsilon$, we stop ray propagation and set the remaining contribution to 0. With the weight ${T_{k}}$ We use MLP predictions to render the color and depth for each ray:
@@
For mapping, we adopt an input keyframe-based approach. Inspired by [1][76][84], we adaptively update the background mask to enhance mapping quality. Unlike traditional dynamic SLAM methods which require complete traversal and completion of all input frames. We only need to perform pixel control on selected keyframes and specific dynamic pixels. We retained static pixels that do not interfere with the tracking process to ensure that our method can correctly distinguish between dynamic and static information without interference.
We obtain the sparse optical flow vectors $V_{s}^{\text{static}}$ for the filtered static feature points mentioned in Sec. 3.1, along with the pixel mask $M_{d}$ in the bounding box. Due to the varying motion speeds of different parts of dynamic objects, additional artifacts are introduced in the form of residual interference. Because the jump sampling strategy filter out low occupancy rate pixels, we can further refine $M_{d}$ by removing low dynamic pixels to reduce its occupancy rate. Based on temporal consistency, we calculate the optical flow vectors $M_{si}^{\text{dynamic}}$ between two consecutive frames for each pixel in $M_{d}$. (In this paper, we use keyframes involved in mapping and their corresponding adjacent frames as reference frames for calculation.) The L1 norm of the difference is obtained as:
$G_{t}=\frac{1}{N}\sum_{i=1}^{N}\left|V_{s}^{\text{static}}-M_{si}^{\text{dynamic}}\right|.$ (8)
We obtain a difference sequence $\{G_{s1},G_{s2},\ldots,G_{sn}\}$. Assuming the sequence follows a normal distribution, we set the threshold as $G_{t}$. When $G_{si}<G_{t}\ \text{{and}}\ 1<G_{si}$, we set the remaining contribution of the corresponding pixel’s depth to 0. To ensure real-time performance and effective removal of low dynamic pixels. If the removed pixels exceed 60% of the total dynamic pixels or the keyframe sampling rate ($n/N$) is less than 60%. We generate a new keyframe from the keyframe list to prevent a decrease in mapping quality due to excessive removal. For the correctness of backend optimization and loop detection, the additionally generated keyframes are not involved in bundle optimization and loop detection.
To supervise the network, we use the color loss $\mathcal{L}_{rgb}$ between the input image $c$ and the rendered image $\hat{c}$ and depth loss. The color and depth rendering loss are set as:
@@
$\mathcal{L}_{rgb}=\frac{1}{N}\sum_{n=1}^{N}\left(\hat{\mathbf{C}}_{s}-\mathbf{C}_{s}\right)^{2},\mathcal{L}_{d}=\frac{1}{\left|R_{s}\right|}\sum_{r\in R_{s}}\left(\hat{\mathbf{D}}_{r}-{\mathbf{D}}_{s}\right)^{2}$ (9)
where $\hat{\mathbf{D}}_{r}$ denotes the filtered ray set with valid depth, and ${\mathbf{D}}_{s}$ corresponds to the filtered pixel depth set. $R_{s}$ is the batch of rays with depth contribution. Specifically, in the monocular mode, we utilize the predicted depth obtained from [75] as supervision. Integrating uncertainty-aware loss, we define the Mapping loss as:
$\mathcal{L}_{M}=\lambda_{r}\mathcal{L}_{\text{rgb }}+\lambda_{d}\mathcal{L}_{\mathrm{D}}.$ (10)
3.3 Tracking and Bundle Adjustment
Camera Tracking. In the presence of ground truth depth, we perform tracking using a system based on [74]. In monocular mode, we utilize the prior generated by [75] to obtain the scale factor $M$. Using the pose matrix $P_{e}$ obtained from the essential matrix, we can express $M$ as:
$d_{c}^{\prime}=K\cdot T\cdot d_{w},\quad M=\frac{1}{N}\sum_{i=1}^{N}\frac{d_{c}^{i}}{d_{p}^{i}}$ (11)
where $d_{c}^{\prime}$ is the feature point in camera coordinate system, $K$ is the camera intrinsic matrix, $T$ is the pose matrix, and $d_{w}$ is the position of the feature point in the world coordinate system. $d_{c}^{i}$ is the predicted depth of the feature point, and $d_{p}^{i}$ is the predicted depth from the network. The corrected pose matrix $P_{c}$ is obtained as $P_{c}=M*P_{e}$.
Loop Detection. In order to achieve effective tracking in large-scale scenes, we introduce loop detection to correct cumulative errors. We retrieve the three most similar keyframes from the keyframe groups using the DBoW2 with a similarity threshold of 0.8. The matching points here refer to the filtered feature points mentioned in Sec3.3. We check for geometric consistency. If at least three out of the 5 co-observed keyframes (already in the map) satisfy the condition of successful matching with the candidate keyframe group, we consider a common area is detected. Once the location recognition is successful, we align the active keyframe $K_{a}$, co-observed keyframe $K_{m}$, and all map points through transformation.
Figure 3: Reconstruction results on the dynamic sequence of TUM RGB-D [49]. The reconstruction results of the RGB-D image stream in high dynamic scenes are shown in the black boxes. The reconstruction results of the RGB-D image stream in low dynamic scenes are shown in the blue boxes. The reconstruction results of the monocular mode in high dynamic sequence are shown in the red boxes. Our method is capable of effectively reconstructing and removing occlusion interference, thus restoring the scale of monocular images.
@@ 3.3 Tracking and Bundle Adjustment
Table 1: The tracking results on 8 dynamic sequences in TUM RGB-D [49](ATE[m]). We also present the average tracking Completion Rate of 6 scenes. The best results are indicated in bold, and the second-best results are indicated with a horizontal line. N/A indicates failure in all 5 measurements and the total number of frames with successful tracking is less than 50% of the total. Our method achieves more stable tracking results compared to traditional methods, while maintaining a running speed of 20fps.
Bundle Adjustment. During pose optimization, we should avoid including dynamic feature points in the optimization process. Therefore, through the filtering process mentioned in Sec. 3.1, we assign weights $w_{i}$ to static points and dynamic points as follows:
$w_{i}=1-{G_{a}}\left(\delta T,T_{\text{th }}\right)$ (12)
where $\delta T$ is the t-statistic of the point and $T_{th}$ is the threshold mentioned in Sec. 3.1. $G_{a}$ is the Gaussian function used to describe the closeness between the point and the background more accurately. We update the weights of each map point based on the corresponding feature points in the current frame. We set the definition of the new joint optimization cost function for the reprojection error as $e_{i}(x)$. To avoid excluding some potential static points, we introduce mean reprojection error $e_{m}(x)$ as a check before BA. It is defined as:
$e_{i}(x)=w_{i}\left|x_{i}-\pi\left(RX_{i}+T_{e}\right)\right|$ (13)
$e_{m}(x)=\frac{1}{N}\sum_{k=1}^{N}\left|x-\pi\left(R_{k}X+T_{ek}\right)\right|$ (14)
where $\pi(\cdot)$ is the projection transformation from the camera coordinate system to the plane coordinate system, and $\rho$ is the robust Huber loss function. $T_{e}$ as the translation vector. We compute the t-value ($T_{em}$) for the average reprojection error to recover deleted feature points, similar to the application of conditional probability for $e_{m}$ as in 3.1. We set the threshold as $T_{eth}$. In particular, we disable the function of eliminating potential feature points in static scenes. In the pose estimation, we jointly optimize the rotation matrix $R$, translation vector $T_{e}$, Mapping of 3D map points to image coordinates $P_{e}$ and the weights $w_{i}$ of each map point. We obtain the minimum reprojection error $L_{rp}$ as follows:
$\min_{\left\{[R\mid T_{e}]\right\},\left\{P_{e}\right\}}L_{rp}=\underset{\left\{[R\mid T_{e}]\right\},\left\{P_{e}\right\}}{\operatorname{argmin}}\sum_{i=0}^{N}\rho\left(e_{i}(x)^{T}\mathbf{\Theta}e_{i}(x)\right)$ (15)
where $\mathbf{\Theta}$ is the information matrix related to the scale of the feature points. We performed BA to optimize the the keyframes and camera poses and the observed map points in these keyframes. The optimized keyframes and mapped points were then passed to NeRF for further processing.
4 Experiment
4.1 Experimental Setup
Implementation Details. Our system is implemented on a system with a NVIDIA RTX 3090ti GPU and an Intel Core i7-12700K CPU operating at 3.60 GHz. We use the YOLOv5 network pre-trained on the COCO dataset[69] as the backbone of our object detection system. Due to the page limit the parameter settings and detailed ablation experiments can be found in the supplementary materials.
Dataset. We evaluated our system on a total of 2 dynamic datasets, 1 challenging dataset with large viewpoints and weak textures, and 3 static datasets. The datasets used for evaluation include TUM RGB-D [49], Bonn [85], OpenLORIS-Scene [77], Replica[48], ScanNet[45], and EuRoC [41]. The TUM RGB-D [49] dataset consists of 6 high-dynamic sequences, 2 low-dynamic sequences, and 3 static sequences. The Bonn [85] dataset includes 6 high-dynamic range sequences and 2 sequences with complex occlusions. The challenging scenes dataset comprises 7 indoor sequences from the OpenLORIS-Scene [77] dataset, demonstrating large viewpoint changes and lighting variations. The static datasets consist of 6 sequences from the ScanNet [45] dataset, 8 sequences from the high-quality virtual scene dataset Replica[48], and 3 Sequences on the EuRoC [41] Dataset. Due to the page limit, we include more experimental results in the supplemental material.
@@ 4.1 Experimental Setup
Table 2: Tracking results on 8 dynamic sequences in the Bonn[85] dataset(ATE[m]). We also present the average tracking Completion Rate of 8 scenes. The best results are indicated in bold, and the second-best results are indicated with a horizontal line. For methods that fail to track, we report the best result obtained before the failure. Our method maintains stable operation at 15FPS while preserving the multi-object tracking capability, aided by TensorRT[79] acceleration.
@@
Figure 4: Reconstruction results on the dynamic sequence of Bonn [85]. Our method is able to reconstruct stably in occluded environments, and compared to traditional methods [54], our approach has advantages in restoring texture details and inferring unknown regions.
Metrics and Baseline. We evaluate the reconstruction quality using Depth L1 (cm), Accuracy (cm), Completion (cm), Completion Ratio (%), and PSNR (dB). For assessing camera tracking quality in dynamic scenes, we use (ATE)RMSE. Our main baselines for reconstruction quality and camera tracking include iMAP[23], NICE-SLAM[42], ESLAM[60], GO-SLAM [31] and Co-SLAM[44]. We utilized the replication code provided by [78] to achieve better performance reproduction for iMAP[23]. For a fair comparison, we introduce classical traditional dynamic SLAM methods for tracking results in dynamic scenes, including LC-CRF SLAM [54], Crowd-SLAM [55], ElasticFusion [4], ORB-SLAM2 [10], and DS-SLAM [20].
4.2 Evaluation in Dynamic Scenes
TUM RGB-D [49] Dynamic Sequence Evaluation. To evaluate our system on TUM RGB-D [49] dynamic sequences, we compare it with LC-CRF SLAM [54], Crowd-SLAM [55], ORB-SLAM2 [10], NICE-SLAM [42], ESLAM [18] and Co-SLAM [44]. Our system achieves effective rendering, effectively avoids occlusion interference caused by high-dynamic human subjects, reasonably fills scene holes, and achieves high-precision geometric detail reconstruction with minimal artifacts, as shown in Figure 4. We have achieved tracking results that are competitive with traditional methods. It is important to highlight that existing neural implicit methods often face challenges in achieving effective tracking in high-dynamic scenes, as shown in Table 2. Due to severe accumulated drift, current neural implicit SLAM methods are unable to complete reliable mapping.
Bonn [85] Dataset Dynamic Sequence Evaluation. Our tracking method demonstrates robustness and achieves competitive tracking results, as demonstrated in Table 3. Compared to the TUM RGB-D [49] dataset, the Bonn [85] dataset exhibits more complex dynamic scenes, severe occlusions, and the presence of multiple moving objects simultaneously. The reconstruction results are shown in Fig. 5. In comparison to traditional dynamic SLAM methods, our approach achieves better results. The current neural implicit systems have shown significant drift or tracking failures in most scenarios.
4.3 Evaluation in Static Scenes
@@
Table 4: Runtime analysis on the ScanNet [45] dataset. The best result is shown in bold, and the second-best result is indicated with an underline. The results of GO-SLAM [31] are from its paper, while the results of other methods are reported based on the average of 6 measurements. Our method demonstrates competitiveness in memory consumption and runtime speed, enabling it to run on lighter-weight platforms, which is a prerequisite for widespread adoption of neural implicit SLAM.
ScanNet [45] Dataset Experimental Results. We assessed the average running speed, memory usage, mapping time, and GPU memory usage (on a 3090ti GPU) across six scenes. The results indicate that our method achieves higher running speed and reduces memory consumption. (Please refer to the supplementary material for detailed tracking results.)
TUM RGB-D [49] Dataset Experimental Results. (including monocular and RGB-D ATE data and reconstruction result images) We tested the performance of monocular and RGB-D modes on this dataset. In three static sequences, our RGB-D mode achieved better tracking results than traditional methods, as shown in Table 4. It changed the disadvantage of neural implicit methods over traditional methods in this dataset. At the same time, our monocular mode also achieved stable tracking results, correctly recovering scale and pose.
Openloris-Scene [77] Dataset Experimental Results. We conducted reconstruction and tracking evaluation on the challenging Openloris-Scene [77] dataset. Our method is significantly better than existing neural implicit methods, and there is no significant drift as in NICE-SLAM [42] and Co-SLAM [44], as shown in Table 5. Thanks to the mature tracking system, our method is able to achieve accurate relocalization in extremely challenging scenes and improve the completion rate of tracking. Even under large viewpoint changes and low texture-low resolution image constraints, we still achieve high-quality reconstruction, which indicates that our 3D reconstruction results benefit from the joint encoding method and adaptive keyframe generation strategy.
[…]
Table 5: Tracking Results on TUM RGB-D [49] dataset. ATE RMSE [cm] (↓) is used as the evaluation metric. We compared neural implicit methods and traditional approaches, and our results achieved superior tracking performance over the traditional methods on the TUM RGB-D [49] dataset. Additionally, our monocular results ensured robustness and accurate scale recovery.

### 09 Mar 2024 v2

@@ 3.3 Dynamic Volume Rendering
For representing 3D scenes in the presence of occlusion, we employ a volumetric rendering process combined with multi-resolution hash encoding [28]. The given position $x$ can be mapped to the respective positions on each resolution grid, and trilinear interpolation is performed on the grid cell corners to obtain concatenated encoding feature vectors $\boldsymbol{\gamma}_{t}(\mathbf{x}_{t},\mathbf{t})\in\mathbb{R}^{c}$ After feature concatenation, we input them into an MLP. More specifically, our SDF network $F_{\Phi_{sdf}}$ is composed of a single MLP, with learnable parameters $\Phi_{sdf}$. It takes the point position $x$ and the corresponding hash encoding $h_{\Phi_{hash}}(\mathbf{x})$ as inputs and predicts the signed distance field (SDF) as:
@@
where $\Omega$ is the sigmoid function. $T_{k}=\prod_{j=1}^{k-1}\left(1-\alpha_{j}\right)$ denotes the cumulative transmission rate, which represents the proportion of light reaching the camera. With the weight ${T_{k}}$ We use MLP predictions to render the color and depth for each ray:
@@
The definitions of color and depth rendering losses are as follows:
@@
$\mathcal{L}_{\text{rgb}}=\frac{1}{N}\sum_{n=1}^{N}\left(\hat{\mathbf{C}}_{n}-\mathbf{C}_{n}\right)^{2},\quad\mathcal{L}_{\text{depth}}=\frac{1}{N}\sum_{n=1}^{N}\left(\hat{\mathbf{D}}_{r}-\mathbf{D}_{r}\right)^{2}$ (16)
where $\hat{\mathbf{D}}_{r}$ represents the set of rays with valid depths, $\mathbf{D}_{r}$ corresponds to the filtered set of pixel depths, and $R_{s}$ is the batch of rays contributing to depth. In monocular mode, supervision is performed using predicted depths from [78]. To effectively reduce the impact of dynamic pixels and improve the accuracy of depth and color predictions, we construct a dynamic rendering loss. Our loss function consists of three parts: the basic loss (including depth loss and color loss), the motion consistency loss, the dynamic area penalty loss ($\mathcal{L}_{\text{pl}}$). The basic loss evaluates the overall accuracy of the prediction, while the additional depth and color penalty items are specifically aimed at dynamic pixel areas to constrain the occlusion impact of dynamic objects. We construct the motion consistency loss ($\mathcal{L}_{\text{m}}$) as follows:
$\mathcal{L}_{\text{m}}=\frac{1}{N_{\text{dynamic}}}\sum_{i\in\text{Dynamic}}\left\|R_{d_{i}}^{\text{pred}}-R_{d_{i}}^{\text{actual}}\right\|^{2}$ (17)
where $R_{d_{i}}^{\text{pred}}$ is the predicted motion distance for the $i$th dynamic pixel, $R_{d_{i}}^{\text{actual}}$ is its actual motion distance, and $N_{\text{dynamic}}$ is the number of dynamic pixels. We construct the dynamic area penalty item ($\mathcal{L}_{\text{pl}}$) as:
$\mathcal{L}_{\text{pld}}=\frac{1}{N_{\text{dynamic}}}\sum_{i\in\text{Dynamic}}\left\|\hat{\mathbf{D}}_{i}-\mathbf{D}_{i}\right\|^{2},\quad\mathcal{L}_{\text{plc}}=\frac{1}{N_{\text{dynamic}}}\sum_{i\in\text{Dynamic}}\left\|\hat{\mathbf{C}}_{i}-\mathbf{C}_{i}\right\|^{2}$ (18)
where $N_{\text{dynamic}}$ is the number of pixels in the dynamic area, $\hat{\mathbf{D}}_{i}$ and $\mathbf{D}_{i}$ respectively represent the predicted and actual depth values. Combining the above losses, we define the rendering loss as:
$\mathcal{L}_{\text{R}}=\lambda_{r}\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{\text{depth}}+\lambda_{\text{dynamic}}(\mathcal{L}_{\text{pld}}+\mathcal{L}_{\text{plc}})+\lambda_{\text{m}}\mathcal{L}_{\text{m}}$ (19)
where $\lambda_{\text{dynamic}}$ and $\lambda_{\text{m}}$ are initialized to 0.001.
4 Experiment
4.1 Experimental Setup
Implementation Details. Our system is implemented on a system with a NVIDIA RTX 3090ti GPU and an Intel Core i7-12700K CPU operating at 3.60 GHz. We use the YOLOv9[84] network pre-trained on the COCO dataset[72] as the backbone of our object detection system. We initialize the optical flow threshold $\delta$ to 5, and the sampling threshold $\theta$ to 500. Due to the page limit the more parameter settings and detailed ablation experiments can be found in the supplementary materials.
Dataset. We evaluated our system on a total of 2 dynamic datasets, 1 challenging dataset with large viewpoints and weak textures, and 3 static datasets. The datasets used for evaluation include TUM RGB-D [52], Bonn [88], OpenLORIS-Scene [80], Replica[51], ScanNet[48], and EuRoC [44]. The TUM RGB-D [52] dataset consists of 6 high-dynamic sequences, 2 low-dynamic sequences, and 3 static sequences. The Bonn [88] dataset includes 6 high-dynamic range sequences and 2 sequences with complex occlusions. The challenging scenes dataset consists of 7 indoor sequences from the OpenLORIS-Scene dataset [80], demonstrating large changes in viewpoint, variations in lighting, and sudden dynamic interferences. The static datasets consist of 6 sequences from the ScanNet [48] dataset, 8 sequences from the high-quality virtual scene dataset Replica[51], and 3 Sequences on the EuRoC [44] Dataset. Due to the page limit, we include more experimental results in the supplemental material.
Metrics and Baseline. We evaluate the reconstruction quality using {Depth L1} (cm), {Accuracy} (cm), {Completion} (cm), {Completion Ratio} (%), and {PSNR} (dB). For assessing camera tracking quality in dynamic scenes, we use {ATE RMSE}. Our main baselines for reconstruction quality and camera tracking include iMAP [26], NICE-SLAM [45], ESLAM [63], GO-SLAM [34], Orbeez-SLAM[20]and Co-SLAM [47]. We utilized the replication code provided by [81] to achieve better performance reproduction for iMAP [26]. For a fair comparison, we introduce classical traditional dynamic SLAM methods for tracking results in dynamic scenes, including LC-CRF SLAM [57], Crowd-SLAM [58], ElasticFusion [4], ORB-SLAM2 [14], and DS-SLAM [23].
4.2 Evaluation in Dynamic Scenes
Figure 4: We compared the reconstruction results of traditional LC-CRF[57] SLAM, ESLAM[21], Orbeez-SLAM[20], and our method (in both RGB and RGB-D modes) on dynamic sequences from TUM RGB-D [52]. We presented results on four TUM dynamic sequences, with two being high-dynamic sequences (involving significant human movement and high speed) and two being low-dynamic sequences (involving minimal human movement and slower speed). Our method demonstrates our ability to preserve potential dynamic objects, eliminate occlusion interference, and reduce the occurrence of artifacts, achieving high-quality reconstruction. Specifically, for a fair comparison, we employed rendering results based on ground truth poses for ESLAM to avoid reconstruction errors caused by tracking inaccuracies.
TUM RGB-D [52] Dynamic Sequence Evaluation. To evaluate our system on TUM RGB-D [52] dynamic sequences, we compare it with LC-CRF SLAM [57], Crowd-SLAM [58], ORB-SLAM2 [14], NICE-SLAM [45], ESLAM [21] and Co-SLAM [47]. Our method achieves effective rendering, effectively avoids occlusion interference caused by high-dynamic human subjects, reasonably fills scene holes, and achieves high-precision geometric detail reconstruction with minimal artifacts. Our method preserves the potential dynamic objects and enhances the integrity of the reconstruction. as shown in Figure 4. We have achieved tracking results that are competitive with traditional methods, as shown in Table 2. Due to severe accumulated drift, current neural implicit SLAM methods are unable to complete reliable mapping.
@@ 4.2 Evaluation in Dynamic Scenes
Table 1: The tracking results on 8 dynamic sequences in TUM RGB-D [52](ATE[m]). We also present the average tracking Completion Rate of 6 scenes. The best results are indicated in bold, and the second-best results are indicated with a horizontal line. N/A indicates failure in all 5 measurements and the total number of frames with successful tracking is less than 50% of the total.
Bonn [88] Dynamic Sequence Evaluation. Our tracking method demonstrates robustness and achieves competitive tracking results, as demonstrated in Table 3. Compared to the TUM RGB-D [52] dataset, the Bonn [88] dataset exhibits more complex dynamic scenes, severe occlusions, and the presence of multiple moving objects simultaneously. The reconstruction results are shown in Fig. 5. In comparison to traditional dynamic SLAM methods, our approach achieves better results. The current neural implicit systems have shown significant drift or tracking failures in most scenarios.
@@
Table 2: Tracking results on 8 dynamic sequences in the Bonn[88] dataset(ATE[m]). We also present the average tracking Completion Rate of 8 scenes. The best results are indicated in bold, and the second-best results are indicated with a horizontal line. For methods that fail to track, we report the best result obtained before the failure.
4.3 Evaluation in Challenging Scenes
Openloris-Scene [80] Dataset Experimental Results. We conducted reconstruction and tracking evaluation on the challenging Openloris-Scene [80] dataset. Our method significantly outperforms existing neural implicit methods and traditional approaches. Benefite from accurate feature point segmentation, and there is no significant drift as in NICE-SLAM [45] and Co-SLAM [47], as shown in Table 5. Our method is capable of tracking in highly challenging scenarios, and improving tracking completion rates.
@@ 4.3 Evaluation in Challenging Scenes
ScanNet [48] Dataset Experimental Results. Our evaluation results are based on the ScanNet[48] dataset(RGB-D), which contains large-scale scenes. We assessed the average running speed, memory usage, mapping time, and GPU memory usage across six scenes. The results indicate that our method achieves higher running speed and reduces memory consumption.
@@
Table 4: Runtime analysis on the ScanNet [48] dataset. The best result is shown in bold, and the second-best result is indicated with an underline. The results of methods are reported based on the average of 6 scenes.